OpenStudy (anonymous):

find the tangents of the curve x^2+ y^2 - 6x + 4y = 0 from (-17, 7). the answers are 2x+3y+13=0 and 34x+129y = 325.. already have the process but i didnt derived to the answer. please help.

OpenStudy (anonymous):

what i did first was i get the slope through dy/dx

OpenStudy (amistre64):

what is your derivative?

OpenStudy (anonymous):

okay wait

OpenStudy (anonymous):

-2x+6/ 2y+4

OpenStudy (amistre64):

there is no need to work it all the way to that step; sinc e y' is what we are after, just input the point values directly. x^2+ y^2 - 6x + 4y = 0 2x+ 2y y' - 6 + 4 y' = 0 2(-17)+ 2(7) y' - 6 + 4 y' = 0 -34 + 14y' - 6 + 4y' = 0 -40 + 18y' = 0 -20 + 9y' = 0 etc

OpenStudy (amistre64):

nothing wrong with working thru all the algebra for y', its just not needed is all

OpenStudy (anonymous):

okay.. so, whats the next step? :)

OpenStudy (amistre64):

our slope is 9/20, so lets see if we get one of those results at least; the next is finding another suitable x,y that fits the bill

OpenStudy (amistre64):

\[y=\frac 9{20}(x+17)+7\] \[20y=9(x+17)+140\] \[20y=9x+294\] \[9x-20y=-294\] now we see if im reading it correctly to begin with :) is -17,7 a solution to the given equation?

OpenStudy (anonymous):

yes yes

OpenStudy (amistre64):

-17,7 is not a solution to: x^2 +y^2 -6x +4y so what it is saying is that that is a point of in the distance that tangents pass thru

OpenStudy (amistre64):



OpenStudy (amistre64):

so we do have to use the general format :)

OpenStudy (amistre64):

2x+ 2y y' - 6 + 4 y' = 0 x+ y y' - 3 + 2y' = 0 y' = (3-x)/(2+y)

OpenStudy (anonymous):

im sorry. i messed up. it is the point on the curve x^2 + y^2 blabla

OpenStudy (amistre64):

for some point(s); (a,b) the follwing system holds \[7 = \frac{3-a}{2+b}(-17-a)+b\] \[a^2+b^2-6a+4b=0\]

OpenStudy (amistre64):

does that make sense? for some tangent slopes coming off of the circle as (a,b), passes thru the point (-17,7) has to satisfy that system

OpenStudy (anonymous):

yes it is..

OpenStudy (amistre64):

\[7 = \frac{3-a}{2+b}(-17-a)+b\] \[7(2+b) = (3-a)(-17-a)+b(2+b)\] \[14+7b = a^2+14a-51+2b+b^2\]hope i saw that last part correctly :) \[65 = a^2+b^2+14a-5b\] you have 2 circle equations that intersect at the "tangent" points needed

OpenStudy (amistre64):

\[a^2+b^2+14a-5b-65=a^2+b^2-6a+4b\] \[14a-5b-65=-6a+4b\] \[20a=9b+65\] \[a=\frac9{20}b+\frac{13}{4}\] now we can produce a quadratic in b, to solve for b

OpenStudy (anonymous):

im following, okay..

OpenStudy (amistre64):

\[a^2+b^2-6a+4b=0\] \[(\frac9{20}b+\frac{13}{4})^2+b^2-6(\frac9{20}b+\frac{13}{4})+4b=0\] solve for b, so that you can solve for a; then use those point to define the lines between (a,b) and (-17,7) I have to get to class for now

OpenStudy (anonymous):

yes! thank you!! so much!! Take care :) have to digest this!! :)